OpenStudy (anonymous):

Solve for x: −2(x + 3) = −2(x + 1) − 4

OpenStudy (ghazi):

-2(x+3)= -2(x+1)-4 -2x-6= -2x-2-4 -2x-6= -2x-6 it can't be solved for x lol

OpenStudy (ghazi):

@davisgurl101 clear?